1.1, -2.2, 3.3, -4.4,...
Simplification.
Formula: sub an=ar(n-1)
(Geometric sequence simplify)

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52802)

.

This sequence is not geometric.

The given formula is irrelevant to given sequence.



Answer by math_tutor2020(3817)


The claim is this sequence is geometric.

To check this claim, divide each term over its previous term.
If we get the same result each time, then the sequence is geometric.

r = term2/term1 = -2.2/1.1 = -2
r = term3/term2 = 3.3/(-2.2) = -1.5
r = term4/term3 = -4.4/3.3 = -1.333333 approximately

We technically could have stopped once reaching -1.5, but it doesn't hurt to check all pairings.
Regardless, this sequence is NOT geometric because there is no common ratio.
We get different result each time we divide adjacent terms.


If the sequence was: 1.1, 2.2, 3.3, 4.4, ...
Then it would be arithmetic because the difference of adjacent terms is the same.
d = term2 - term1 = 2.2 - 1.1 = 1.1
d = term3 - term2 = 3.3 - 2.2 = 1.1
d = term4 - term3 = 4.4 - 3.3 = 1.1